OpenStudy (anonymous):

Evaluate arctan(tan(5pi/6))? Exact answer

OpenStudy (nikvist):

\[\arctan: (-\infty,+\infty)\rightarrow [-\pi/2,\pi/2]\]\[\arctan{(\tan{(5\pi/6)})}=5\pi/6-\pi=-\pi/6\]